# VIP: Versatile Image Outpainting Empowered by Multimodal Large Language Model

### v2

@@ 4.1. Dataset
Laion2B-en (Schuhmann et al., 2021, 2022) The dataset contains about 2B image-text data pairs. We apply the data generation strategy to generate 224 pieces of data with the proposed text prompt format.
@@
Scenery (Yang et al., 2019) The Scenery dataset consists of many images of natural landscapes and has about 5,000 images for training and 1,000 images for testing.
@@
Building Facades (Gao et al., 2023) The images in this dataset are focused on urban buildings. There are about 16,000 images in the training set and 1,500 images in the testing set.
@@
WikiArt (Tan et al., 2016) WikiArt is a large dataset of paintings by 195 artists, which contains about 45,500 training images and 19,500 testing images. It has different degrees of realism and stylization.
Figure 4. The visualization comparison of the outpainting results generated by different methods. Our proposed approach surpasses other well-behaved methods in the concerned box region.
[…]
A woman and a girl in white dresses sit in an open car.
4.2. Implementation Details
@@
In this section, the implementation details are described briefly. Based on the pretrained stable-diffusion-inpainting model, we perform two training processes. The one, denoted specific training (SFT), is fine-tuning the model on Scenery, Building, and Wikiart dataset for comparison of the SoTA results with the existing methods in the proposed unconditional type, ’Center:; Surrounding:’, respectively. As existing methods do, the input size is set to $192\times 192$. Unlike the traditional hundreds of epochs for training, the fine-tuning iteration is set to only 30000 with a mini-batch 4. The original images from the datasets are first resized to the set size as ground-truth images. The contents of the region with a size $128\times 128$ in the center are preserved, and the surrounding region is masked off. To show the generalization ability of the proposed method and achieve customized results, the other, namely generalized training (GLT), is fine-tuning the model on the data generated by MLLM from Laion2B-en dataset. With an input size of $256\times 256$, the fine-tuning iteration is set to 10000 with a mini-batch 4. A common optimizer, Adam (Kingma and Ba, 2014), is selected with a learning rate of 5e-6. In the inference stage, the center size and the extrapolation ratio are set to the same as the corresponding training process.
@@
For evaluation, the commonly used Frechet Inception distance (FID) (Heusel et al., 2017) and Inception Score (IS) (Salimans et al., 2016) are chosen as the indicators to show the quantitative performance.
4.3. Compared with SoTA methods
@@
In this section, we compare the outpainting results obtained by the proposed method with those obtained by existing SoTA methods, and the superiority of the proposed method is proven. For quantitative comparison, we select the methods, SRN (Wang et al., 2019), NSIPO (Yang et al., 2019), IOH (Van Hoorick, 2019) , U-Transformer (Gao et al., 2023), Neural ODE (Gao et al., 2024), QueryOTR(Yao et al., 2022) and PQDiff (Zhang et al., 2024b). The statistical data of FID and IS indicators are recorded in Table 1.
@@
As shown in the third and fourth lines from the bottom, performing generalized training (GLT) can achieve a comparable result with the well-behaved methods, indicating that our proposed method has a strong generalization ability and can extend to the outpainting tasks on various image types with more realistic and consistent results. In addition, after performing specific training (SFT) on the three datasets, the results far exceed those of existing methods, as shown in the last two lines. The FID indicator achieves the best level at 10.717, 17.775, and 3.010 on the Scenery, Building, and Wikiart datasets, respectively. Meanwhile, the IS indicator also achieves excellent and advanced results. Especially it can be seen that we do not need to perform fine-tuning on the wikiart dataset and can achieve the SoTA FID results just based on the fine-tuning model on the generated data from the Laion-2B dataset.
Table 2. Results of using each component in the proposed method
@@
        CS         16.827         4.031
        CS + CTS         14.883         4.123
In addition, the leading results are achieved by just fine-tuning only one-digit epochs instead of fine-tuning the model for a hundred epochs as existing models, which greatly saves training costs based on the strong generation ability of the pretrained model. Meanwhile, it can be found that whether or not replacing the center content of the generated results with the original center content, our results both exceed the existing best FID result. This means our method can sense the spatial position relation and generate the content of the center and surrounding regions precisely. Moreover, the FID boost effect is not as obvious as the previous method after replacing the center content, which also proves our method can generate whole images with consistent and continuous content.
We also compare the visualization results with other methods. For a fair comparison, we select the same image cases as research (Gao et al., 2024) as shown in Fig 11. Inside the selected box regions, it can be found that the texture details are more clearly visible in the surrounding regions. For example, in the first line, the edge of the dune is smoother and clearer than that of other methods. The outpainting performance has been greatly improved in terms of overall consistency and continuity.
4.4. Ablation Study
In this section, we analyze the effectiveness of each component in the proposed learning method through some ablation experiments on the test set of the Scenery dataset (Yang et al., 2019). We set the outpainting ratio to 2.25x, like most works. The size of the input center region is 128x128, and the size of the output extrapolation image is 192x192. Both qualitative and quantitative results are provided below.
Figure 5. The visualization comparison of the ablation results of each component in the proposed method.
[…]
A woman and a girl in white dresses sit in an open car.
Table 3. Results of using different unconditional paradigms
@@ 4.4. Ablation Study
Components of The Proposed Framework. To show the contribution of each component, the quantitative results are demonstrated in Table 4. It is evident that the contribution of each component to the final success of our proposed method is important. The first line shows results obtained by directly using the pretrained stable-diffusion-inpainting model. The unsatisfactory results demonstrate that the outpainting task is fundamentally different from the inpainting task, and the model optimized for the inpainting task can not be directly used in the outpainting task to obtain a satisfactory result. After fine-tuning the model on the proposed image-text data pair with space information, ’Center:; Surrounding’ (CS), the FID indicator has been greatly improved, as shown in the second line. This means the distribution of extrapolated images is more similar to that of the ground truth images. The IS indicator can also maintain a comparable level. Meanwhile, the effectiveness of CTS cross-attention is also verified by further using it, and the results are shown in the third line. The FID and IS indicators are both improved compared to those obtained by only using the proposed text prompt format.
Figure 6. The visualization comparison of using different unconditional paradigms.
@@
We further analyze the effectiveness of each component through the qualitative results as shown in Fig. 5. As we can see, the baseline can not behave well in the outpainting task without fine-tuning. With the proposed prompt format, the outpainting results are obviously improved. Moreover, adding the CTS cross-attention further improves the texture details and reduces abnormal generations.
@@
vs. Traditional Unconditional Prompt Format. In this part, we compare the fine-tuning results obtained by the traditional unconditional prompt format, prompt = ’, and the proposed unconditional prompt format, ’Center:; Surrounding:’. The results are recorded in Table 3. It can be found that our proposed format can achieve better results in the FID indicator with a negligible decrease in the IS indicator. The visualization of the results is also demonstrated in Fig. 6. It can be found the our method is better than the traditional unconditional format regarding reality and consistency. The performance of the unconditional outpainting results has been improved after utilizing the proposed method.
Figure 7. The visualization comparison of the portrait results.
Figure 8. The visualization comparison of the customized outpainting results with different text prompts in diverse scenes.
4.5. Portrait Results
In this section, we demonstrate the outpainting results for a more difficult scene, portrait generation. The visualization results on the WikiArt (Tan et al., 2016) dataset are provided in Fig. 7. It can be found that the extrapolated parts are seamless with the original portrait, and the body parts are continuous.
4.6. Customized Results
@@
Thanks to the pretrained stable-diffusion-inpainting model, customized outpainting can be more easily achieved based on the strong generation ability with text prompt conditions after fine-tuning. In this section, the fine-tuned model by performing the generalized training (GLT) is utilized for the customized outpainting. Then we select four samples randomly to provide the corresponding visualization results. Based on the proposed method, we can achieve customized outpainting results according to the proposed text prompt condition, as shown in Fig 8 (a).
@@
For regular outpainting, it can be found that the generation results change correspondingly after adding different surrounding keywords. For example, if we set the keyword of the surrounding region to ’shirt’, ’cap’, ’sign’, and ’box’, the generated results correspondingly change to an accurate object. Therefore, precise conditional customized outpainting can be achieved through our proposed method. In addition, we can also utilize the MLLM to imagine the surrounding content and output more abundant keywords instead of an artificial mode. Then, the various outpainting results can be obtained without a tedious human design process.
In addition, as shown in Fig. 8 (b), we also perform the customized outpainting with an irregularly shaped mask to show the generalization ability of our approach. The noteworthy object in the center region often has an irregular shape. Therefore, we utilize the Grounded SAM (Ren et al., 2024) to extract the main objects in the center region and mask the surrounding region. Then we obtain the masked image and the corresponding mask to perform the irregular outpainting task.
@@
For example, it can be found that in the first case, according to elements in the provided prompt, we can continuously add the surrounding content of the airplane with ’desert’, ’runway’, ’mountain’, and ’sky’, and the corresponding content will appear in the generated result. Moreover, as we continue to add keywords, the generated result maintains good consistency and continuity on the basis of the original result. Therefore, our model can not extrapolate the region of common square shape like the existing methods but also extrapolate the region of irregular shape. Our model can adapt to various outpainting scenarios and obtain satisfactory results.
5. Conclusion
@@
In this paper, we propose a novel framework namely VIP to complete the image outpainting task with Multimodal Large Language model. The method can introduce text prompts with the space information of center and surrounding regions to perform precise outpainting. With the help of the Multimodal Large Language model, we design a new text prompt format called ’Center:;Surrounding:’ (CS) to assign the different object components to specific regions. Then, the proposed novel Center-Total-Surrounding (CTS) cross-attention can further enhance the correlation between each image region and the corresponding text part through an adaptive fusion of multiple cross-attention operations. Therefore, the model can be aware of the exclusive characters of different regions and generate consistent results across the whole image. We carry out extensive experiments on widely-used datasets. Leading quantitative and qualitative results are obtained in both unconditional and conditional outpainting tasks. The effectiveness of our method is verified. In future work, we intend to continue to explore a more effective text prompt format for customized outpainting results.
Appendix A Overview of Appendix
In this supplementation material, we provide additional details that were omitted in the main manuscript owing to the limited space. As for implementation details, we first give more description of the experiment environment and settings and the usage of the Center-Total-Surrounding (CTS) cross-attention. Then, more experiments are conducted to show the effectiveness of the proposed approach, containing a comparison with an open-source outpainting method and an analysis on more different hyperparameter settings. Finally, the visualization comparison with other SoTA methods on more examples is demonstrated. Moreover, various outpainting results obtained by our proposed method are provided below.
Appendix B More Implementation Details
B.1. Training
In this paper, we utilize the pretrained stable diffusion inpainting with an input channel of the unet as 9, which has been fine-tuned explicitly for inpainting. In the fine-tuning process, the whole parameters of the unet are optimized, and other parts like the text-encoder and the vae are frozen. Our sampler utilizes the DDIM (Ho et al., 2020) sampler with 50 sampling timesteps. In the training stage, there is only a traditional loss function of calculating the mean-square error (MSE) between the predicted noise and added noise. The experiments are carried out on the NVIDIA V100 with Pytorch 1.13.
B.2. Center-Total-Surrounding Cross-Attention
For Center-Total-Surrounding (CTS) Cross-Attention, we introduce the module in the downsampling, middle, and upsampling processes, where the original cross-attention module exists. After calculating the resnet layer, the three cross-attention modules are calculated and the corresponding results are adaptively fused for the subsequent forward process.
B.3. The examples of The Generated Training Data
In this section, we will demonstrate some examples of the training data generated by MLLM. The selected image-text pair samples are visualized in Fig. 9. In the future, we will release the whole generated data.
Figure 9. Qualitative samples generated by MLLM.
[…]
A woman and a girl in white dresses sit in an open car.
It can be seen that the content keywords of the center and surrounding regions are effectively split in the four generated samples. With our proposed approach, we can make full use of data in this format for better outpainting performance.
Appendix C More Experiments
Figure 10. Customized results compared with the open-source outpainting method, SD-infinity Outpainting.
[…]
A woman and a girl in white dresses sit in an open car.
C.1. Comparison of Customized Results
In the main manuscript, we have made a comparison with the leading methods reported in the published paper. To further show the advantage of our work, we further compare it with a unpublished but open-sourced method  [^1] https://github.com/lkwq007/stablediffusion-infinity, namely SD-infinity Outpainting here, which is also based on the pretrained stable diffusion of the traditional prompt. With the same setting of the sampling parameters, the visualization results are demonstrated in Fig 10.
It can be found that after using the proposed approach, the best outpainting results are obtained, and the proposed text prompt paradigm with just some keywords is relatively simple compared to the traditional prompt paradigm with a complete sentence. In the first line, the dividing line between sand and road is clearer and more rational. In the second line, the cap generated by our approach has a more normal shape and a more continuous color. In the third line, the waist area is more realistic, and there are fewer tattoos on the arms In the last line, the generated lane obtained by the open-source outpainting repository is more disorganized.
Table 4. Results of using different parameter formats
        Method         FID $\downarrow$         IS $\uparrow$
        baseline         15.421         4.131
        RIF         15.170         4.216
        CF ($a=0.5$)         15,270         4.224
        LF         14.883         4.123
C.2. Analysis of Different Hyperparameter Setting
In this section, we analyze the influence of setting the hyperparameter $a$ used in the CTS cross-attention module. Specifically, we propose three variations for comparison with our used strategy: the random initialization format (RIF), the constant format (CF), and the used learnable format (LF). The corresponding results on scenery are recorded in Table 4.
It can be found that no matter which format we use, the FID results obtained are always better than those obtained by the baseline, and the IS indicator maintains an excellent level. This indicates that our method is insensitive to the parameter format. The best FID results are achieved by utilizing the learnable format compared to the other two formats. It is understandable that directly setting a constant or random initialization will not adapt to the fine-tuning process and obtain better results. Moreover, using the random initialization format is better than the constant format, which means that using the constant format needs to perform extensive experiments to explore the best value, which is very consuming. Therefore, the learnable format is a reasonable and effective strategy.
Appendix D More Results
D.1. Traditional Unconditional Results
Figure 11. The demonstration of more samples compared to the other methods on three datasets.
[…]
A woman and a girl in white dresses sit in an open car.
In this section, we provide the visualization of other samples used in research (Gao et al., 2024). The unconditional outpainting results on the Scenery, Building, and Wikiart datasets are demonstrated in Fig. 11. As we can see, the texture details are more pronounced around the edges, and the contents are more continuous and not as abnormal as in other methods. For example, the extrapolated ship in the first line has a more normal shape, and the building in the fourth line has more structural information.
Appendix E More Results
E.1. Traditional Unconditional Results
Figure 12. The demonstration of more outpainting results on the three datasets.
[…]
A woman and a girl in white dresses sit in an open car.
In addition, we provide more outpainting results generated by our approach on the three datasets in Fig. 12. Left, middle, and right are the Scenery, Building Facades, and WikiArt dataset, respectively. For input, only the part in the black box area of the GT image is reserved, and the other parts are extrapolated. It can be found that the extrapolated part is consistent with the input part, and the outpainting performance is satisfactory on various samples.
E.2. Customized Conditional Results
In this part, more conditional customized results obtained on samples from the Laion-2B dataset are also shown in Fig. 13. It can be seen that with various prompts, we can obtain visually plausible customized outpainting results.
Figure 13. The demonstration of more customized outpainting results.
[…]
A woman and a girl in white dresses sit in an open car.

### v3

@@ 4.1 Dataset
Laion2B-en [35, 34] The dataset contains about 2B image-text data pairs. We apply the data generation strategy to generate 224 pieces of data with the proposed text prompt format.
@@
Scenery [45] The Scenery dataset consists of many images of natural landscapes and has about 5,000 images for training and 1,000 images for testing.
@@
Building Facades [10] The images in this dataset are focused on urban buildings. There are about 16,000 images in the training set and 1,500 images in the testing set.
@@
WikiArt [37] WikiArt is a large dataset of paintings by 195 artists, which contains about 45,500 training images and 19,500 testing images. It has different degrees of realism and stylization.
MSCOCO [18] The MSCOCO dataset features over 330,000 images that depict a wide variety of everyday scenes and objects, including people, animals, and items. Each image is accompanied by five distinct natural language descriptions, providing diverse perspectives on the visual content. In this case, the test set consisting of approximately 40,000 images.
4.2 Implementation Details
@@
In this section, the implementation details are described briefly. Based on the pre-trained stable-diffusion-inpainting model, we perform two training processes. The one, denoted specific training (SFT), is fine-tuning the model on Scenery, Building, and Wikiart dataset for comparison of the SoTA results with the existing methods in the proposed unconditional type, ’Center:; Surrounding:’, respectively. As existing methods do, the input size is set to $192\times 192$. Unlike the traditional hundreds of epochs for training, the fine-tuning iteration is set to only 30000 with a mini-batch 4. The original images from the datasets are first resized to the set size as ground-truth images. The contents of the region with a size $128\times 128$ in the center are preserved, and the surrounding region is masked off. To show the generalization ability of the proposed method and achieve customized results, the other, namely generalized training (GLT), is fine-tuning the model on the data generated by MLLM from the Laion2B-en dataset. With an input size of $256\times 256$, the fine-tuning iteration is set to 10000 with a mini-batch 4. A common optimizer, Adam [17], is selected with a learning rate of 5e-6. In the inference stage, the center size and the extrapolation ratio are set to the same as the corresponding training process.
Table 2: Quantitative results under different mask types on MSCOCO
Method Random Mask Outpainting Mask
FID $\downarrow$ LPIPS $\downarrow$ FID $\downarrow$ LPIPS $\downarrow$
LaMa [36] 27.21 0.3137 9.52 0.1283
SD1.5-inpainting [28] 10.29 0.3879 5.35 0.1512
SD3-inpainting [8] 8.94 0.3465 -- --
SD3-inpainting Turbo [33] 9.44 0.3416 -- --
VIP 9.24 0.2005 1.57 0.1092
Figure 4: Visualization comparison bewteen different methods on the Scenery dataset. VIP can surpass other SoTA methods in the concerned box region.
@@
For evaluation, the commonly used Frechet Inception distance (FID) [12] and Inception Score (IS) [32] are chosen as the indicators of quantitative performance.
4.3 Comparisons with State-of-The-Art
@@
Text-Unconditional Generation. In this section, we compare the outpainting results obtained by the proposed method with those obtained by existing SoTA methods, and the superiority of the proposed method is proven. For quantitative comparison, we select the methods, SRN [41], NSIPO [45], IOH [39] , U-Transformer [10], Neural ODE [11], QueryOTR[46] and PQDiff [51]. The statistical data of FID and IS indicators are recorded in Table 1.
@@
As shown in the third and fourth lines from the bottom, performing generalized training (GLT) can achieve a comparable result with the well-behaved methods, indicating that our proposed method has a strong generalization ability and can extend to the outpainting tasks on various image types with more realistic and consistent results. In addition, after performing specific training (SFT) on the three datasets, the results far exceed those of existing methods, as shown in the last two lines. The FID indicator achieves the best level at 10.717, 17.775, and 3.010 on the Scenery, Building, and Wikiart datasets, respectively. Meanwhile, the IS indicator also achieves excellent and advanced results. Especially it can be seen that we do not need to perform fine-tuning on the Wikiart dataset and can accomplish the SoTA FID results just based on the fine-tuning model on the generated data from the Laion-2B dataset.
In addition, the leading results are achieved by just fine-tuning only one-digit epochs instead of fine-tuning the model for a hundred epochs as existing models, which greatly saves training costs based on the strong generation ability of the pre-trained model. Meanwhile, it can be found that whether or not replacing the center content of the generated results with the original center content, our results both exceed the existing best FID result. This means our method can sense the spatial position relation and generate the content of the center and surrounding regions precisely. Moreover, the FID boost effect is not as obvious as the previous method after replacing the center content, which also proves our method can generate whole images with consistent and continuous content.
We also provide visualization comparison between various methods. As shown in Fig 11, with our method, the generated content is more clearly visible and consistent in the surrounding regions inside the selected box regions. For example, in the second line, the edge of the dune is smoother and clearer than that of other methods. The outpainting performance has been greatly improved in terms of overall consistency and continuity.
Figure 5: Visualization comparison under different outpainting mask types.
Text-Conditional Generation. To further demonstrate the superiority of the proposed method in text-conditional generation, as done in SD3-Turbo [33], the quantitative comparison on MSCOCO under random mask types is provided in Table 2. In addition, we also provide results only under the outpainting mask type. It can be seen that the performance of VIP far exceeds that of SD1.5-inpainting, especially in outpainting, and reaches a level comparable to SD3 which uses a complex model structure. In addition, the visualization comparison is also demonstrated in Fig. 5. For different outpainting positions, the performance of our method can exceed that of LaMa and SD1.5 inpainting, in terms of the continuity and reasonableness.
4.4 Ablation Study
In this section, we analyze the effectiveness of each component in VIP through various ablation experiments on the test set of the Scenery dataset [45]. We set the outpainting ratio to 2.25x, like most works. The size of the input center region is 128x128, and the size of the output extrapolation image is 192x192. Both qualitative and quantitative results are provided below.
Table 3: Effectiveness of each component
@@ 4.4 Ablation Study
w/ CS 16.827 4.031
w/ CS + C-T-S 14.883 4.123
Table 4: Comparison of using different unconditional paradigms
@@
Figure 6: Visualization comparison of using each component.
Figure 7: Visualization comparison with different unconditional paradigms.
@@
Effectiveness of Each Component. To show the contribution of each component in our method, quantitative results are demonstrated in Table 6. It is evident that the contribution of each component to the final success of our proposed method is important. The first line shows results obtained by directly using the pre-trained stable-diffusion-inpainting model. The unsatisfactory results demonstrate that the outpainting task is fundamentally different from the inpainting task, and the model optimized for inpainting can not be directly used in the outpainting task for satisfactory results. After fine-tuning the model on the proposed image-text data pair with space information, ’Center:; Surrounding’ (CS), the FID indicator has been greatly improved, as shown in the second line. The IS indicator can also maintain a comparable level. Meanwhile, the effectiveness of C-T-S decoupled control mechanism is also verified by further using it, and the results are shown in the third line. The FID and IS indicators are both improved compared to those obtained by only using the proposed text prompt format. This means the introduction of space information helps the model perceive the unique properties of original and extrapolated regions.
@@
We further analyze the effectiveness of each component through the qualitative results as shown in Fig. 7. As we can see, the baseline can not behave well in the outpainting task without fine-tuning. After fine-tuning with the proposed prompt format, the outpainting results are obviously improved. Moreover, adding the C-T-S decoupled control mechanism further improves the texture details and reduces abnormal generations.
@@
vs. Traditional Unconditional Prompt Format. In this part, we compare the fine-tuning results obtained by the traditional unconditional prompt format, prompt = ’, and the proposed unconditional prompt format, ’Center:; Surrounding:’. The results are recorded in Table 4. It can be found that our proposed format can achieve better results in the FID indicator with a negligible decrease in the IS indicator. The visualization of the results is also demonstrated in Fig. 7. It can be found that our method is better than the traditional unconditional format regarding reality and consistency. The performance of the unconditional outpainting results has been improved after utilizing the proposed method.
Table 5: Results of using different captions for training
Method Clip Score
w/ general caption 28.2435
w/ CS 29.4943
w/ CS + C-T-S 30.3588
Analysis of Text-Image Consistency. In this section, we demonstrate the superiority of VIP in terms of the precise text condition by evaluating the clip score as shown in Table 5. As shown in the second line, just using CS, the proposed text caption with space information, can improve the text-image consistency. In addition, after using the C-T-S decoupled control mechanism, a better clip score is obtained by enhancing control of text over corresponding image regions.
4.5 Various Customized Results
@@
Customized outpainting can be more easily achieved based on text prompt conditions after using the proposed method. In this section, the fine-tuned model by performing the generalized training (GLT) is utilized to achieve customized outpainting. Then we select two samples randomly for visualization. Based on the proposed method, we can achieve customized outpainting results according to the proposed text prompt condition.
@@
For regular outpainting as shown in Fig 8 (a), it can be found that generated results change after adding different surrounding keywords. If we set the keyword of the surrounding region to ’shirt’, ’cap’, ’sign’, and ’box’, generated results correspondingly change to the accurate object. Therefore, precise conditional customized outpainting is achieved through our proposed method. In addition, we can also utilize the MLLM to imagine the surrounding content and output more abundant keywords instead of an artificial mode.
@@
In addition, as demonstrated in Fig. 8 (b), we also perform the customized outpainting with an irregularly shaped mask to show the generalization ability. The noteworthy object in the center region often has an irregular shape. Therefore, we utilize the Grounded SAM [27] to extract foreground objects and mask the surrounding region. Then we obtain the masked image and the corresponding mask to perform the irregular outpainting task. According to elements in the provided prompt, we continuously add the surrounding content with ’ocean’, ’road’, ’smooth’, and ’fence’, and the corresponding content will appear in the generated result. Moreover, as we continue to add keywords, the generated result maintains good consistency and continuity based on the original result. Therefore, our model can not extrapolate the region of common square shape like the existing methods but also extrapolate the region of irregular shape. Our model can adapt to various outpainting scenarios and obtain satisfactory results.
Figure 8: Visualization comparison of customized outpainting results with different text prompts in diverse scenes.
5 Conclusion
@@
In this paper, we propose a novel framework namely VIP to complete the image outpainting task with the Multimodal Large Language model. The method can introduce text prompts with the space information of the center and surrounding regions to perform precise outpainting. With the help of the Multimodal Large Language model, we design a new text prompt format called ’Center:; Surrounding:’ (CS) to assign different object components to specific regions. Then, the proposed Center-Total-Surrounding (C-T-S) decoupled control mechanism can further enhance the correlation between each image region and the corresponding text part through an adaptive fusion of multiple cross-attention operations. Therefore, the model can be aware of the exclusive characters of different regions and generate consistent results across the whole image. We carry out extensive experiments on widely-used datasets. Leading quantitative and qualitative results are obtained in both unconditional and conditional outpainting tasks. The effectiveness of our method is verified. In future work, we intend to continue to explore a more effective text prompt format for customized outpainting results.